# Global weak solutions for the inverse mean curvature flow in the Heisenberg group

### v2

@@ 2. Basic notations and preliminary results
    $\|g\|_{\varepsilon}\leq\|g\|$ for any $g\in\mathbb{H}^{n}$, whence $B_{r}(\bar{g})\subseteq\mathcal{B}^{\varepsilon}_{r}(\bar{g})$ for any $\bar{g}\in\mathbb{H}^{n}$ and for any $r>0$. Moreover, $\|g\|_{\varepsilon}\geq\frac{1}{2}\|g\|$ whenever $g\in\mathbb{H}^{n}$ satisfies $\|g\|_{\varepsilon}\geq 2\varepsilon$.
@@
    If $0<\rho_{1}<\rho_{2}\leq 2\rho_{1}$ and $h:[0,\infty)\to[0,1]$ is a Lipschitz and piecewhise linear function, with $\{h\equiv 1\}=[0,\rho_{1}]$, $\{h\equiv 0\}=[\frac{\rho_{1}+\rho_{2}}{2},\infty)$, and $h$ is linear otherwise, then for any $\bar{g}\in\mathbb{H}^{n}$ the function $\zeta(g)=h(\|\bar{g}^{-1}*g\|_{\varepsilon})$ is piecewise $C^{1}$ on $\mathbb{H}^{n}$ with $\zeta\equiv 1$ on $\mathcal{B}^{\varepsilon}_{\rho_{1}}(\bar{g})$, $spt\,\zeta\subset\mathcal{B}^{\varepsilon}_{\rho_{2}}(\bar{g})$ and
@@
Clearly we may assume $\hat{g}\neq\bar{g}$, otherwise the claim holds trivially for $\rho\leq r$ with $C=1$ and in view of (2.2) for $\rho\in(r,2r]$. For fixed $\varepsilon>0$ the space $(\mathbb{H}^{n},d_{\varepsilon})$ is a locally compact Riemannian manifold which is complete as a metric space. It follows from the Hopf-Rinow Theorem that for any $\hat{g}\in B^{\varepsilon}_{r}(\bar{g})$, $\hat{g}\neq\bar{g}$, there exists a (strictly positive) constant speed geodesic $\Gamma:[0,1]\to\mathbb{H}^{n}$ such that $\Gamma(0)=\bar{g}$, $\Gamma(1)=\hat{g}$ and $d_{\varepsilon}(\Gamma(t),\Gamma(t^{\prime}))=|t-t^{\prime}|d_{\varepsilon}(\bar{g},\hat{g})$ for any $t,t^{\prime}\in[0,1]$. Note that the last identity obviously yields $\Gamma(t)\in B^{\varepsilon}_{r}(\bar{g})$ for any $t\in[0,1]$.
Suppose first that $\rho\in[\frac{1}{4}r,2r]$, then we also have $\Gamma(t)\in B^{\varepsilon}_{\rho}(\hat{g})$ for any $t\in[\frac{3}{4},1]$. In addition, $B^{\varepsilon}_{\rho}(\hat{g})\subset B^{\varepsilon}_{3r}(\bar{g})$ by triangle inequality. Since $\tilde{g}:=\Gamma(7/8)\in B^{\varepsilon}_{\rho}(\hat{g})\cap B^{\varepsilon}_{r}(\bar{g})$ and $d_{\varepsilon}(\tilde{g},\bar{g})=7d_{\varepsilon}(\tilde{g},\hat{g})=\frac{7}{8}d_{\varepsilon}(\bar{g},\hat{g})$, then $B^{\varepsilon}_{r/{16}}(\tilde{g})\subset B^{\varepsilon}_{\rho}(\hat{g})\cap B^{\varepsilon}_{r}(\bar{g})$. Indeed for any $g\in B^{\varepsilon}_{r/{16}}(\tilde{g})$ we have $d_{\varepsilon}(g,\bar{g})\leq d_{\varepsilon}(g,\tilde{g})+d_{\varepsilon}(\tilde{g},\bar{g})<r/16+\frac{7}{8}r<r$ and, analogously, $d_{\varepsilon}(g,\hat{g})\leq d_{\varepsilon}(g,\tilde{g})+d_{\varepsilon}(\tilde{g},\hat{g})<r/16+r/8<\rho$.
@@
Finally, suppose instead that $40<\rho<r/$ 4. Clearly we may assume $d_{\varepsilon}(\bar{g},\hat{g})\geq\frac{3}{4}r$, because otherwise $B_{\rho}(\hat{g})\subset B_{r}(\bar{g})$ and (2.8) trivially holds with $C=1$. Then by continuity there exists $t_{1}\in(0,1)$ such that $d_{\varepsilon}(\Gamma(t_{1}),\hat{g})=\rho$ and $d_{\varepsilon}(\Gamma(t),\hat{g})<\rho$ for any $t\in(t_{1},1]$. Now we set $t_{2}=\frac{1}{2}(t_{1}+1)$ and $\tilde{g}=\Gamma(t_{2})\in B^{\varepsilon}_{\rho}(\hat{g})$, so that we clearly have $B^{\varepsilon}_{\rho}(\hat{g})\subset B^{\varepsilon}_{2\rho}(\tilde{g})$. Note that $d_{\varepsilon}(\tilde{g},\bar{g})=d_{\varepsilon}(\bar{g},\hat{g})-d_{\varepsilon}(\tilde{g},\hat{g})<r-\rho/2$ and therefore $B^{\varepsilon}_{\rho/{4}}(\tilde{g})\subset B^{\varepsilon}_{\rho}(\hat{g})\cap B^{\varepsilon}_{r}(\bar{g})$. Indeed for any $g\in B^{\varepsilon}_{\rho/4}(\tilde{g})$ we have $d_{\varepsilon}(g,\bar{g})\leq d_{\varepsilon}(g,\tilde{g})+d_{\varepsilon}(\tilde{g},\bar{g})<\rho/4+r-\rho/2<r$ and, analogously, $d_{\varepsilon}(g,\hat{g})\leq d_{\varepsilon}(g,\tilde{g})+d_{\varepsilon}(\tilde{g},\hat{g})<\rho/4+\rho/2<\rho$.
@@ 3. Construction of barriers
In this section, in the range $1<p<Q$ we provide a family of barriers whch will be used for establishing the boundary gradient estimate for $p$-capacitary potentials. These functions are negative powers of the Korányi norm obtained as suitable deformation of those giving singular $p$-harmonic functions w.r.to the sub-Riemannian metric. As a result of a careful choice of the exponent, they will be singular subsolutions for the sub-Riemannian $p$-Laplace operator in the whole $\mathbb{H}^{n}$ and for the Riemannian $p$-Laplace operator in the complement of any fixed Korányi ball.
We first introduce a polynomial function that will be used frequently in the computation. For any $g=(x,y,t)\in\mathbb{H}^{n}$ we set
@@
Consequently, for $\varepsilon\in(0,1]$ we also have
(3.14) $|\nabla_{\varepsilon}N|_{\varepsilon}^{2}=|\nabla_{0}N|_{0}^{2}+\varepsilon^{2}(TN)^{2}=16|z|^{2}N+64\varepsilon^{2}16t^{2}\leq 16N^{3/2}+64N\,.$
@@
For $p>1$ and $K>0$ let $\alpha=-\frac{K}{p-1}<0$ and $\Phi^{(g_{0},R_{0})}_{\alpha}$ as in (3.30). Then $\Phi^{(g_{0},R_{0})}_{\alpha}=1$ on $\partial B_{R_{0}}(g_{0})$ and there exist $C_{0}>0$ depending only on $R_{0}$ such that for any $\varepsilon\in(0,1]$ we have
(3.31) $\||\nabla_{\varepsilon}\Phi^{(g_{0},R_{0})}_{\alpha}|_{\varepsilon}\|_{L^{\infty}(\partial B_{R_{0}}(g_{0}))}\leq\frac{C_{0}K}{p-1}\,.$
Moreover, for the function $\tilde{\Phi}^{(g_{0},R_{0})}_{\alpha}=\min\{1,\Phi^{(g_{0},R_{0})}_{\alpha}\}$ we have $\tilde{\Phi}^{(g_{0},R_{0})}_{\alpha}\in\dot{W}_{\varepsilon}^{1,p}(\mathbb{H}^{n})$ whenever $K>\frac{(Q-p)(p-1)}{4p}$ and the norms are uniformly bounded for $\varepsilon\in(0,1]$.
@@
Up to translations we may assume $g_{0}=0$. Next for $\Phi=\Phi^{(0,R_{0})}_{\alpha}$ inequality (3.14) yields
(3.32) $|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}=\left|\nabla_{\varepsilon}\left(\frac{N}{R_{0}^{4}}\right)^{\alpha}\right|_{\varepsilon}^{2}=\alpha^{2}\Phi^{2}\frac{|\nabla_{\varepsilon}N|_{\varepsilon}^{2}}{N^{2}}\leq\alpha^{2}\Phi^{2}(64N^{-1}+16N^{-1/2})\,.$
Evaluating on $\partial B_{R_{0}}(0)$ gives the conclusion with $C_{0}=\frac{4}{R_{0}^{2}}\sqrt{4+R_{0}^{2}}$.
Concerning integrability, assuming up to translations $g_{0}=0$ and denoting $A_{j}=\{2^{j}R_{0}<\|\cdot\|\leq 2^{j+1}R_{0}\}$, for any $j\geq 0$, by change of variables we have $\mathcal{H}^{Q}(A_{j})\simeq(2^{Q})^{j}$ and in turn
$\int_{\|\cdot\|>R_{0}}\Phi_{\alpha}^{p^{*}}=R_{0}^{-4\alpha p^{*}}\sum_{j=0}^{\infty}\int_{A_{j}}\left(N^{\alpha}\right)^{p^{*}}\simeq\sum_{j=0}^{\infty}(2^{4\alpha p^{*}})^{j}\mathcal{H}^{Q}(A_{j})\simeq\sum_{j=0}^{\infty}(2^{4\alpha p^{*}+Q})^{j}\,.$
As a consequence $\tilde{\Phi}^{(g_{0},R_{0})}_{\alpha}\in L^{p^{*}}$ if and only if $04\alpha p^{*}+Q<$ 0, which holds for $K>0$ precisely as in the statement above.
@@
$\int_{\|\cdot\|>R_{0}}|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{p}\lesssim\sum_{j=0}^{\infty}\int_{A_{j}}\left(N^{\alpha}\right)^{p}N^{-p/4}\simeq\sum_{j=0}^{\infty}(2^{(4\alpha-1)p})^{j}\mathcal{H}^{Q}(A_{j})\simeq\sum_{j=0}^{\infty}(2^{(4\alpha-1)p+Q})^{j}\,,$
@@
(3.42) $-N^{\gamma}\geq-R_{0}^{4\gamma},\quad\textrm{for every }\gamma<0.$
@@
(3.45) $(III)$ $=\varepsilon^{4}(p-1)(T\Phi)^{2}(TT\Phi)=\varepsilon^{4}(p-1)(32\alpha N^{\alpha-1}t)^{2}\left[32\alpha N^{\alpha-2}(N+(\alpha-1)32t^{2})\right]$
$=\varepsilon^{4}(32\alpha)^{3}(p-1)N^{3\alpha-4}t^{2}\left[|z|^{4}+\underbrace{16(2\alpha-1)t^{2}}_{\leq 0}\right],$
@@
Note that in order to construct subsolutions in a Riemannian case one has to adjust the exponent comparing to the one in the sub-Riemannian case treated in Lemma 3.1. However, the difference of the exponents corresponding to the two choices can be made smaller and smaller as $\varepsilon$ gets smaller or $R_{0}>0$ gets larger. We will exploit this fact in the proof of Proposition 5.1. On the other hand, the construction of Riemannian supersolutions via neagative powers of the Korányi norm seems to be impossible. Indeed, considering (3.41), (3.43) and (3.45) restricted to the vertical axis (i.e. for $|z|=0$) clearly gives the opposite of the desired sign. A similar behaviour will be faced even considering the $\varepsilon$-gauge defined in (2.6), as they agree with the Korányi norm for $|t|\geq\varepsilon$.
4. Two uniform estimates in mustiscale geometries.
@@
In this section we present two results concerning $p$-harmonic functions on the Riemannian Heisenberg group $(\mathbb{H}^{n},d_{\varepsilon})$ which for fixed $p\in(1,Q)$ hold uniformly w.r.to the parameter $\varepsilon\in(0,1]$, namely, an $L^{p}-L^{\infty}$ estimates for their gradient in Proposition 4.1 and a local Harnack inequality in Proposition 4.2. As it will be clear in the next sections, the these two results will be essential in Proposition 6.1 in order to establish decay properties at infinity for Riemannian $p$-capacitary potentials in exterior domains and in turn to prove the gradient bounds announced in Theorem 1.2 and Corollary 1.3.
@@
Note that the $L^{p}-L^{\infty}$ estimates for their gradient discussed here is the counterpart of the one obtained in [37] for the horizontal gradient of $p$-harmonic functions on the sub-Riemannian Heisenberg group. The result presented here is essentially contained in [10], although not explicitely stated there. Indeed, in [10] the key estimate concerning a weighted Caccioppoli inequality for the gradient is obtained and here we fill the missing details to derive the claim through uniform local Sobolev inequality from [19] and Moser iteration. Similarly, the uniform local Harnack inequality discussed here is the Riemannian counterpart of the one proved in [11] for the sub-Riemannian case (see also [9] for analogous results concerning the Riemannian approximation for parabolic equations modeled on the $p$-Laplacian for $p\geq 2$). Here, following closely the argument in [11] for the sub-Riemannian case, we detail how to obtain the Harnack inequality for $p$-harmonic functions in the Riemannian case uniformly on $\varepsilon\in(0,1]$, relying on the main results in [19] (see also [9]), i.e., the uniform doubling property and the uniform Poincaré inequalities reviewed in Section 2 in combination with Moser iteration and John-Nirenberg inequality.
@@
Since we were not able to find a reference in the literature and although the arguments are essentially standard and well-known to the experts, for all the three results in the sequel we sketch the proofs mainly for the reader’s convenience, so that uniformity w.r.to $\varepsilon\in(0,1]$ will be transparent and the dependence on $p$ will be made explicit when needed.
@@
For fixed $\varepsilon\in(0,1]$ and $B^{\varepsilon}_{2r}(\bar{g})\subseteq\Omega$ we consider the auxiliary domain $\Omega^{\prime}:=\mathcal{B}^{\varepsilon}_{2\bar{C}^{-1}r}(\bar{g})\subseteq B^{\varepsilon}_{2r}(\bar{g})$, i.e., a ball associated to the $\varepsilon$-gauge $\|\cdot\|_{\varepsilon}$ discussed in Lemma 2.2 for a constant $\bar{C}>1$ depending only on $Q$ obtained there. Thus, chosing $\theta=2\bar{C}^{2}>1$ and setting for brevity $\bar{r}=(2\bar{C})^{-1}r$, Lemma 2.2-(3) yields $B^{\varepsilon}_{\theta^{-1}r}(\bar{g})\subseteq\mathcal{B}^{\varepsilon}_{\bar{r}}(\bar{g})\subseteq\mathcal{B}^{\varepsilon}_{2\bar{r}}(\bar{g})\subseteq B^{\varepsilon}_{r}(\bar{g})$ and the lemma is proved once we show that
@@
(4.3) $\left\{\begin{array}[]{ll}{\rm div}_{\varepsilon}\left(\left(\sigma+|\nabla_{\varepsilon}w|^{2}_{\varepsilon}\right)^{(p-2)/2}\nabla_{\varepsilon}w\right)=0&\textrm{in }\Omega^{\prime},\\ w=v&\textrm{on }\partial\Omega^{\prime}\,,\end{array}\right.$
@@
(4.6) $\int_{\Omega^{\prime}}\eta^{2}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)^{\frac{p-2+\beta}{2}}|\nabla^{2}_{\varepsilon}w^{\sigma}|^{2}\leq C\beta^{10}\left(\||\nabla_{\varepsilon}\eta|_{\varepsilon}\|_{L^{\infty}}^{2}+\|\eta T\eta\|_{L^{\infty}}\right)\int_{spt\,\eta}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)^{\frac{p+\beta}{2}}\,,$
where $C>0$ depends only on $p$ and $Q$ and $|\nabla^{2}_{\varepsilon}w^{\sigma}|^{2}$ is the sum of the squared second derivatives w.r. to the orthonormal frame $\{X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n},T_{\varepsilon}\}$.
Now notice that $\Omega^{\prime}=\mathcal{B}_{4\bar{r}}^{\varepsilon}(\bar{g})$ and chosing $\bar{r}\leq\rho_{1}<\rho_{2}\leq 2\bar{r}\leq 2\rho_{1}\leq 4\bar{r}$ we have $\mathcal{B}_{\rho_{1}}^{\varepsilon}(\bar{g})\subset\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})\subseteq\Omega^{\prime}$, so that by a simple limiting argument inequality (4.6) still holds for $\eta=\zeta$, the Lipschitz cut-off function given in Lemma 2.2-(5). Thus, combining (4.6) with the elementary inequality $|\nabla_{\varepsilon}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)|^{2}_{\varepsilon}\leq 4\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)|\nabla^{2}_{\varepsilon}w^{\sigma}|^{2}$ we obtain for any $\beta\geq 2$
@@
In order to conclude the proof it remains to show how to infer (4.5) from (4.10), i.e., how to decrease the exponent from $p+2$ to $p$ on the r.h.s. of (4.10) (while inflating slightly the corresponding ball), and to do this we adapt the interpolation argument in [27, Lemma 3.38]. First, using (4.8) with $\beta=2$ for any $\bar{r}\leq\rho_{1}<\rho_{2}\leq 2\bar{r}\leq 2\rho_{1}\leq 4\bar{r}$, so that $\lambda=\rho_{1}/\rho_{2}\in[\frac{\bar{r}}{\rho_{2}},1)\subseteq[\frac{1}{2},1)$, we deduce that the following inequality holds
@@
then in view of Hölder’s inequalty, (4.11) and $1<p<Q$ we have for fixed $\rho_{2}\in[\bar{r},2\bar{r}]$
@@
Finally, chosing $\nu=\frac{1}{2\widetilde{C}}$ and letting $\delta\to 0$ in (4.15) we get $\Phi(\rho_{2};p)\leq 2c(\nu,\mu,p,Q)\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\rho_{2}}(\bar{g})|}\int_{\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})}\Psi_{\sigma}^{p}\right)^{\frac{1}{p}}$, which together with (4.13) for $\rho_{1}=\frac{3}{2}\bar{r}$, $\rho_{2}=2\bar{r}$ and $\lambda=\frac{\rho_{1}}{\rho_{2}}=\frac{3}{4}>\frac{1}{2}$ give

### v3

@@ 2. Basic notations and preliminary results
    $\|g\|_{\varepsilon}\leq\|g\|$ for any $g\in\mathbb{H}^{n}$, whence $B_{r}(\bar{g})\subseteq\mathcal{B}^{\varepsilon}_{r}(\bar{g})$ for any $\bar{g}\in\mathbb{H}^{n}$ and for any $r>0$. Moreover, $\|g\|_{\varepsilon}\geq\frac{1}{2}\|g\|$ whenever $g\in\mathbb{H}^{n}$ satisfies $\|g\|_{\varepsilon}\geq\varepsilon$.
@@
    If $0<\rho_{1}<\rho_{2}\leq 2\rho_{1}$ and $h:[0,\infty)\to[0,1]$ is a Lipschitz and piecewise linear function, with $\{h\equiv 1\}=[0,\rho_{1}]$, $\{h\equiv 0\}=[\frac{\rho_{1}+\rho_{2}}{2},\infty)$, and $h$ is linear otherwise, then for any $\bar{g}\in\mathbb{H}^{n}$ the function $\zeta(g)=h(\|\bar{g}^{-1}*g\|_{\varepsilon})$ is piecewise $C^{1}$ on $\mathbb{H}^{n}$ with $\zeta\equiv 1$ on $\mathcal{B}^{\varepsilon}_{\rho_{1}}(\bar{g})$, $spt\,\zeta\subset\mathcal{B}^{\varepsilon}_{\rho_{2}}(\bar{g})$ and
@@
[…]
@@
[…]
@@ 3. Construction of barriers
In this section, in the range $1<p<Q$ we provide a family of barriers which will be used for establishing the boundary gradient estimate for $p$-capacitary potentials. These functions are negative powers of the Korányi norm obtained as suitable deformation of those giving singular $p$-harmonic functions w.r.to the sub-Riemannian metric. As a result of a careful choice of the exponent, they will be singular subsolutions for the sub-Riemannian $p$-Laplace operator in the whole $\mathbb{H}^{n}$ and for the Riemannian $p$-Laplace operator in the complement of any fixed Korányi ball.
We first introduce a polynomial function that will be used frequently in the computations. For any $g=(x,y,t)\in\mathbb{H}^{n}$ we set
@@
Consequently, for $\varepsilon\in(0,1]$ we also have $(T_{\varepsilon}N)^{2}\leq 64\varepsilon^{2}N$ and
(3.14) $|\nabla_{\varepsilon}N|_{\varepsilon}^{2}=|\nabla_{0}N|_{0}^{2}+\varepsilon^{2}(TN)^{2}=16|z|^{2}N+64\varepsilon^{2}16t^{2}\leq 16N^{3/2}+64\varepsilon^{2}N\,.$
@@
For $p>1$ and $K>0$ let $\alpha=-\frac{K}{p-1}<0$ and $\Phi^{(g_{0},R_{0})}_{\alpha}$ as in (3.30). Then $\Phi^{(g_{0},R_{0})}_{\alpha}=1$ on $\partial B_{R_{0}}(g_{0})$ and for any $\varepsilon\in(0,1]$ we have
(3.31) $\||\nabla_{0}\Phi^{(g_{0},R_{0})}_{\alpha}|_{0}\|_{L^{\infty}(\partial B_{R_{0}}(g_{0}))}\leq\frac{4K}{R_{0}(p-1)}\,\qquad\hbox{and}\quad\|T\Phi^{(g_{0},R_{0})}_{\alpha}\|_{L^{\infty}(\partial B_{R_{0}}(g_{0}))}\leq\frac{8K}{R_{0}^{2}(p-1)}.$
Moreover, for $\tilde{\Phi}^{(g_{0},R_{0})}_{\alpha}=\min\{1,\Phi^{(g_{0},R_{0})}_{\alpha}\}$ we have $\tilde{\Phi}^{(g_{0},R_{0})}_{\alpha}\in\dot{W}_{\varepsilon}^{1,p}(\mathbb{H}^{n})$ whenever $K>\frac{(Q-p)(p-1)}{4p}$ and the Sobolev norms are uniformly bounded for $\varepsilon\in(0,1]$.
@@
Up to translations we may assume $g_{0}=0$. Next for $\Phi=\Phi^{(0,R_{0})}_{\alpha}$ inequality (3.14) together with (3.17) -(3.19) yield
(3.32) $|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{2}=\left|\nabla_{\varepsilon}\left(\frac{N}{R_{0}^{4}}\right)^{\alpha}\right|_{\varepsilon}^{2}=\alpha^{2}\Phi^{2}\frac{|\nabla_{0}N|_{0}^{2}+\varepsilon^{2}(TN)^{2}}{N^{2}}\leq\alpha^{2}\Phi^{2}(16N^{-1/2}+64\varepsilon^{2}N^{-1})\,.$
Evaluating on $\partial B_{R_{0}}(0)$ inequality (3.32) easily yields (3.31).
Concerning integrability, assuming up to translations $g_{0}=0$ and denoting $A_{j}=\{2^{j}R_{0}<\|\cdot\|\leq 2^{j+1}R_{0}\}$, for any $j\geq 0$, by change of variables we have $|A_{j}|\simeq(2^{Q})^{j}$ and in turn
$\int_{\|\cdot\|>R_{0}}\Phi_{\alpha}^{p^{*}}=R_{0}^{-4\alpha p^{*}}\sum_{j=0}^{\infty}\int_{A_{j}}\left(N^{\alpha}\right)^{p^{*}}\simeq\sum_{j=0}^{\infty}(2^{4\alpha p^{*}})^{j}|A_{j}|\simeq\sum_{j=0}^{\infty}(2^{4\alpha p^{*}+Q})^{j}\,.$
As a consequence $\tilde{\Phi}^{(g_{0},R_{0})}_{\alpha}\in L^{p^{*}}$ if and only if $4\alpha p^{*}+Q<0$, which holds for $K>0$ precisely as in the statement above.
@@
$\int_{\|\cdot\|>R_{0}}|\nabla_{\varepsilon}\Phi|_{\varepsilon}^{p}\lesssim\sum_{j=0}^{\infty}\int_{A_{j}}\left(N^{\alpha}\right)^{p}N^{-p/4}\simeq\sum_{j=0}^{\infty}(2^{(4\alpha-1)p})^{j}|A_{j}|\simeq\sum_{j=0}^{\infty}(2^{(4\alpha-1)p+Q})^{j}\,$,
@@
(3.42) $N^{\gamma}\geq-R_{0}^{4\gamma},\quad\textrm{for every }\gamma<0.$
@@
[…]
@@
Note that in order to construct subsolutions in a Riemannian case one has to adjust the exponent comparing to the one in the sub-Riemannian case treated in Lemma 3.1. However, the difference of the exponents corresponding to the two choices can be made smaller and smaller as $\varepsilon$ gets smaller or $R_{0}>0$ gets larger. We will exploit this fact in the proof of Proposition 5.1. On the other hand, the construction of Riemannian supersolutions via negative powers of the Korányi norm seems to be impossible. Indeed, considering (3.41), (3.43) and (3.45) restricted to the vertical axis (i.e. for $|z|=0$) clearly gives the opposite of the desired sign. A similar behaviour will be faced even considering the $\varepsilon$-gauge defined in (2.8), as they agree with the Korányi norm for $|t|\geq\varepsilon$.
4. Two uniform estimates in multiscale geometries.
@@
In this section we present two results concerning $p$-harmonic functions on the Riemannian Heisenberg group $(\mathbb{H}^{n},d_{\varepsilon})$ which for fixed $p\in(1,Q)$ hold uniformly w.r.to the parameter $\varepsilon\in(0,1]$, namely, an $L^{p}-L^{\infty}$ estimates for their gradient in Proposition 4.1 and a local Harnack inequality in Proposition 4.2. As it will be clear in the next sections, the these two results will be essential in Proposition 8.1 in order to establish decay properties at infinity for the gradient of Riemannian $p$-capacitary potentials in exterior domains and in turn to prove the gradient bounds announced in Theorem 1.2 and Corollary 1.3.
@@
Note that the $L^{p}-L^{\infty}$ estimates for their gradient discussed here is the counterpart of the one obtained in [39] for the horizontal gradient of $p$-harmonic functions on the sub-Riemannian Heisenberg group. The result presented here is essentially contained in [10], although not explicitly stated there. Indeed, in [10] the key estimate concerning a weighted Caccioppoli inequality for the gradient is obtained and here we fill the missing details to derive the claim through uniform local Sobolev inequality from [20] and Moser iteration. Similarly, the uniform local Harnack inequality discussed here is the Riemannian counterpart of the one proved in [11] for the sub-Riemannian case (see also [9] for analogous results concerning the Riemannian approximation for parabolic equations modeled on the $p$-Laplacian for $p\geq 2$). Here, following closely the argument in [11] for the sub-Riemannian case, we detail how to obtain the Harnack inequality for $p$-harmonic functions in the Riemannian case uniformly on $\varepsilon\in(0,1]$ and with a quantitative dependence on $p$ , relying on the main results in [20] (see also [9]), i.e., the uniform doubling property and the uniform Poincaré and Sobolev inequalities reviewed in Section 2 in combination with Moser iteration and John-Nirenberg inequality.
@@
Since we were not able to find a reference in the literature and although the arguments are essentially standard and well-known to the experts, for both the aforementioned results we sketch in the sequel the proofs mainly for the reader’s convenience, so that uniformity w.r.to $\varepsilon\in(0,1]$ will be transparent and the dependence on $p$ will be made explicit when needed.
@@
For fixed $\varepsilon\in(0,1]$ and $B^{\varepsilon}_{2r}(\bar{g})\subseteq\Omega$ we consider the auxiliary domain $\Omega^{\prime}:=\mathcal{B}^{\varepsilon}_{2\bar{C}^{-1}r}(\bar{g})\subseteq B^{\varepsilon}_{2r}(\bar{g})$, i.e., a ball associated to the $\varepsilon$-gauge $\|\cdot\|_{\varepsilon}$ discussed in Lemma 2.2 for a constant $\bar{C}>1$ depending only on $Q$ obtained there. Thus, choosing $\theta=2\bar{C}^{2}>1$ and setting for brevity $\bar{r}=(2\bar{C})^{-1}r$, Lemma 2.2-(3) yields $B^{\varepsilon}_{\theta^{-1}r}(\bar{g})\subseteq\mathcal{B}^{\varepsilon}_{\bar{r}}(\bar{g})\subseteq\mathcal{B}^{\varepsilon}_{2\bar{r}}(\bar{g})\subseteq B^{\varepsilon}_{r}(\bar{g})$ and the lemma is proved once we show that
@@
[…]
@@
(4.6) $\int_{\Omega^{\prime}}\eta^{2}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)^{\frac{p-2+\beta}{2}}|\overline{D}^{2}_{\varepsilon}w^{\sigma}|^{2}\leq C\beta^{10}\left(\||\nabla_{\varepsilon}\eta|_{\varepsilon}\|_{L^{\infty}}^{2}+\|\eta T\eta\|_{L^{\infty}}\right)\int_{spt\,\eta}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)^{\frac{p+\beta}{2}}\,$,
where $C>0$ depends only on $p$ and $Q$ and $|\overline{D}^{2}_{\varepsilon}w^{\sigma}|^{2}$ is the sum of the squared second derivatives w.r. to the orthonormal frame $\{X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n},T_{\varepsilon}\}$.
Now notice that $\Omega^{\prime}=\mathcal{B}_{4\bar{r}}^{\varepsilon}(\bar{g})$ and choosing $\bar{r}\leq\rho_{1}<\rho_{2}\leq 2\bar{r}\leq 2\rho_{1}\leq 4\bar{r}$ we have $\mathcal{B}_{\rho_{1}}^{\varepsilon}(\bar{g})\subset\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})\subseteq\Omega^{\prime}$, so that by a simple limiting argument inequality (4.6) still holds for $\eta=\zeta$, the Lipschitz cut-off function given in Lemma 2.2-(5). Thus, combining (4.6) with the elementary inequality $|\nabla_{\varepsilon}\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)|^{2}_{\varepsilon}\leq 4\left(\sigma+|\nabla_{\varepsilon}w^{\sigma}|^{2}_{\varepsilon}\right)|\overline{D}^{2}_{\varepsilon}w^{\sigma}|^{2}$ we obtain for any $\beta\geq 2$
@@
In order to conclude the proof it remains to show how to infer (4.5) from (4.10), i.e., how to decrease the exponent from $p+2$ to $p$ on the r.h.s. of (4.10) (while inflating slightly the corresponding ball), and to do this we adapt the interpolation argument in [28, Lemma 3.38]. First, using (4.8) with $\beta=2$ for any $\bar{r}\leq\rho_{1}<\rho_{2}\leq 2\bar{r}\leq 2\rho_{1}\leq 4\bar{r}$, so that $\lambda=\rho_{1}/\rho_{2}\in[\frac{\bar{r}}{\rho_{2}},1)\subseteq[\frac{1}{2},1)$, we deduce that the following inequality holds
@@
then in view of Hölder inequality, (4.11) and $1<p<Q$ we have for fixed $\rho_{2}\in[\bar{r},2\bar{r}]$
@@
Finally, choosing $\nu=\frac{1}{2\widetilde{C}}$ and letting $\delta\to 0$ in (4.15) we get $\Phi(\rho_{2};p)\leq 2c(\nu,\mu,p,Q)\left(\frac{1}{|\mathcal{B}^{\varepsilon}_{\rho_{2}}(\bar{g})|}\int_{\mathcal{B}_{\rho_{2}}^{\varepsilon}(\bar{g})}\Psi_{\sigma}^{p}\right)^{\frac{1}{p}}$, which together with (4.13) for $\rho_{1}=\frac{3}{2}\bar{r}$, $\rho_{2}=2\bar{r}$ and $\lambda=\frac{\rho_{1}}{\rho_{2}}=\frac{3}{4}>\frac{1}{2}$ give